Che Required a. The following table contains missing data which should be completed. Use the letters-NA to indicate that no amount is related to the alpha character shown in the model. Use the letters FA to represent financing activity, IA to represent investing activities, OA to represent operating activities, and NC to represent net change. Balance Sheet Income Statement -LiabilitiesStockholders Equity Flows CommonEarnings LandPayable Beg NA NA| 0l + NA NA NA NA 200,000 FA NAI 350,000 350,000| + NAI + NAI+ NAI+ 350,000+ NA+ NAI+ NA (500.000) NA9500095,000 OA (65,000)(65,000) OA 95,000 5 (65,000) NAI NA- NA (12,000) NAII NAII NA 18,000 65,000 30,000 68,000 NC < Prev 2 3ds-4 5 of 12 Hi Next >
Answer #1

THE FRAME SHOP (TFS)

Horizontal Financial Statement Model for year 1

Balance Sheet

Income Statement

Statement of Cash Flows

Assets

=

Liabilities

+

Shareholder's Equity

Revenue

-

Expense

=

Net Income

Cash

+

land

=

Notes payable

+

Common Stock

+

Retained Earnings

-

=

Beg.

$ 0.00  

+

$ 0.00

=

$ 0.00

+

$ 0.00

+

$ 0.00

$ 0.00

-

$ 0.00

=

$ 0.00

N/A

1

$    200,000.00

+

N/A

=

N/A

+

$ 200,000.00

+

N/A

N/A

-

N/A

=

N/A

$    200,000.00

FA

2

$    350,000.00

+

N/A

=

$ 350,000.00

+

N/A

+

N/A

N/A

-

N/A

=

N/A

$    350,000.00

FA

3

$ (500,000.00)

+

$ 500,000.00

=

N/A

+

N/A

+

N/A

N/A

-

N/A

=

N/A

$ (500,000.00)

IA

4

$      95,000.00

+

N/A

=

N/A

+

N/A

+

$   95,000.00

$ 95,000.00

-

N/A

=

$   95,000.00

$      95,000.00

OA

5

$    (65,000.00)

+

N/A

=

N/A

+

N/A

+

$ (65,000.00)

N/A

-

$ 65,000.00

=

$ (65,000.00)

$    (65,000.00)

Oa

6

$    (12,000.00)

+

N/A

=

N/A

+

N/A

+

$ (12,000.00)

N/A

-

N/A

=

N/A

$    (12,000.00)

FA

End.

$      68,000.00

+

$ 500,000.00

=

$ 350,000.00

+

$ 200,000.00

+

$   18,000.00

$ 95,000.00

-

$ 65,000.00

=

$   30,000.00

$      68,000.00

NC

Notes

  1. Cash paid in entry no 6 represents Dividends distributed during the period. It is a financing activity.
  2. Ending balance of cash account will be equal to balance of cash in Cash flow statement column.
  3. Land is purchased for $500000 as can be seen in cash flow column cash flow is Investing activity. There is only one asset which is land so land is purchased for $500000.
  4. Common stock is purchased for 200000 as can be seen from cash flow statement column. Cash flow states 200000 as financing activity.
  5. All others are also balancing figures.

Feel free for queries related to this question in the comment section.